(2x-5)^2+4(2x-5)+3=0
Answer by stanbon(75887)

(2x-5)^2+4(2x-5)+3=0
-------
Let 2x-5 = w
===================
w^2+4w+3 = 0
Factor:
(w+3)(w+1) = 0
w = -3 or w = -1
--------------------
Solve for "x":
2x-5 = -3
2x = 2
x = 1
------
OR
---
2x-5 = -1
2x = 4
x = 2
-------
Final Solution: x = 1 or x = 2
================================
Cheers,
Stan H.
==============